# New Results on the Polyak Stepsize: Tight Convergence Analysis and Universal Function Classes

### v2

@@ 4 Universal Convergence of $\gamma$-PolyakGD Across Function Classes
In this section, we demonstrate that $\gamma$-PolyakGD is universal across various function classes, following the terminology “universal” introduced in nesterov2015universal. We first establish convergence under Hölder smoothness and growth conditions, and then provide several extensions, such as relaxations of convexity and the stochastic Polyak stepsize.
@@ 4.1 Convergence results under Hölder smoothness and growth
The above theorem demonstrates that $\gamma$-PolyakGD adapts to both Hölder smoothness and Hölder growth conditions. Specifically, letting $r\rightarrow\infty$, the above result reduces to $\mathcal{O}(1/K^{(\nu+1)/2})$, which matches the rate achieved by the universal gradient method in nesterov2015universal. This result is shown to be tight in Theorem 3.3. On the other hand, setting $\nu=0$, the above result reduces to $\mathcal{O}(1/K^{r/(2(r-1))})$, which matches the lower bound (nemirovskii1985optimal). This rate is also achieved by restart-based optimization methods (roulet2017sharpness; yang2018rsg; xu2019accelerating; d2021acceleration; renegar2022simple; grimmer2023general).
@@ 4.2 Extensions
in Equation (14). This inequality corresponds to a weaker condition known as star-convexity (see Definition 4.1 in nesterov2018lectures, for example). Therefore, the convexity assumption can naturally be relaxed to star-convexity. For example, when $f(x)\coloneqq h(g(x))$ is the composition between an outer scalar function $h:\mathbb{R}\rightarrow\mathbb{R}$ and a function $g:\mathbb{R}^{n}\rightarrow\mathbb{R}$, $f$ can be shown to be star-convex under mild conditions of $h$ and $g$.
@@
This structure is common in machine learning optimization tasks (chen2021solving; jiang2022optimal), where the function $f$ itself may be nonconvex and nonsmooth. Star-convexity suffices to ensure that Theorem 4.1 remains valid. The only modification is that we need to use the subgradient of $f$ in $\gamma$-PolyakGD (Algorithm 1).
@@
Note that Theorem 4.1 is not applicable to the case $\gamma=2$. To establish a convergence guarantee for this case, we introduce the following inequality. This inequality is an extension of inequality (2.1.10) [^4] This inequality is also known as cocoercivity in the literature (see, e.g., malitsky2024adaptive; latafat2024adaptive). in nesterov2018lectures to Hölder smoothness.
@@
We now consider a more general upper curvature bound than Hölder smoothness, known as the global curvature bound (nesterov2025universal; doikov2025universal). This bound is intrinsic as it does not require the objective function $f$ to belong to any specific function class; it is defined by the following quantity:
@@
Given $\hat{\mu}_{f}$, its inverse function $s_{f}$, known as the complexity gauge, corresponds to a stationary criterion for $f$ (see Definition 1 in nesterov2025universal). Specifically, it satisfies $\hat{\mu}_{f}(s_{f}(t))=t$ for all $t\geq 0$. The following theorem shows that $\gamma$-PolyakGD also adapts to this global curvature bound, matching the result of the universal primal gradient method in nesterov2025universal.
@@
To conclude this section, we extend $\gamma$-PolyakGD to stochastic Polyak stepsize (loizou2021stochastic). Consider the following stochastic optimization problem
@@ Appendix A Technical Lemmas
According to Theorem 6.3 in rodomanov2020smoothness, the function $p(x)=\|x\|^{1+\nu}$ is $(\nu,2^{1-\nu}(\nu+1))$-Hölder smooth. Note that $\nabla p_{A}(x)=\nabla(p(Ax))=A\nabla p(Ax)$. Hence, for any $x,y\in\mathbb{R}^{n}$, we have
@@ Appendix B Missing Proofs in Section 3
First, observe that the Huber loss is a quadratic function with condition number $1$ when $\|x\|\leq 1/(2K+1)$. By Theorem 1 in hazan2019revisiting, 1-PolyakGD converges to $x^{\star}$ in one step once the iterate enters this region. Hence, it suffices to consider an initial point $x^{0}$ satisfying $\|x^{0}\|>1/(2K+1)$. We now show that after a single iteration, the iterate $x^{1}$ generated by $1$-PolyakGD satisfies $\|x^{1}\|\leq 1/(2K+1)$. Specifically, the stepsize is given by
@@ Appendix C Missing Proofs in Section 4
Let $\hat{\sigma}_{f}(t)\coloneqq\int_{0}^{t}\frac{\hat{\mu}(\tau)}{\tau}\operatorname{d}\tau$. By Theorem 1 in nesterov2025universal, it holds that
@@
where $(2\hat{\sigma}_{f})^{*}$ is the Fenchel conjugate function of $2\hat{\sigma}_{f}$ (see Definition 3.22 in drusvyatskiy2020convex). Consequently, by the Fenchel-Young inequality, it follows $\zeta\|\nabla f(x)\|\leq 2\hat{\sigma}_{f}(\zeta)+(2\hat{\sigma}_{f})^{*}(\|\nabla f(x)\|)$. Note that $f(x)-f^{\star}\geq(2\hat{\sigma}_{f})^{*}(\|\nabla f(x)\|)$. We derive
@@
Again, recall that $\zeta$ is arbitrary. This proves Equation (23). Since the function $f$ is convex, by Lemma 4 in nesterov2025universal, it holds that $\hat{\sigma}_{f}(t)\leq\hat{\mu}_{f}(t)$ for any $t\in[0,\infty)$. From Equation (23), we obtain

### v3

@@ 4 Universal Convergence of $\gamma$-PolyakGD Across Function Classes
In this section, we demonstrate that $\gamma$-PolyakGD is universal across various function classes, following the terminology “universal” introduced in Nesterov (2015). We first establish convergence under Hölder smoothness and growth conditions, and then provide several extensions, such as relaxations of convexity and the stochastic Polyak stepsize.
@@ 4.1 Convergence results under Hölder smoothness and growth
The above theorem demonstrates that $\gamma$-PolyakGD adapts to both Hölder smoothness and Hölder growth conditions. Specifically, letting $r\rightarrow\infty$, the above result reduces to $\mathcal{O}(1/K^{(\nu+1)/2})$ [^4] Orabona and D’Orazio (2025) established a more general result for the stochastic setting under Hölder smoothness with arbitrary surrogate functions; see Theorem 7 therein., which matches the rate achieved by the universal gradient method in Nesterov (2015). This result is shown to be tight in Theorem 3.3. On the other hand, setting $\nu=0$, the above result reduces to $\mathcal{O}(1/K^{r/(2(r-1))})$, which matches the lower bound (Nemirovskii and Nesterov, 1985). This rate is also achieved by restart-based optimization methods (Roulet and d’Aspremont, 2017; Yang and Lin, 2018; Xu, 2019; d’Aspremont et al., 2021; Renegar and Grimmer, 2022; Grimmer, 2023).
@@ 4.2 Extensions
in Equation (14). This inequality corresponds to a weaker condition known as star-convexity (see Definition 4.1 in Nesterov and others (2018), for example). Therefore, the convexity assumption can naturally be relaxed to star-convexity. For example, when $f(x)\coloneqq h(g(x))$ is the composition between an outer scalar function $h:\mathbb{R}\rightarrow\mathbb{R}$ and a function $g:\mathbb{R}^{n}\rightarrow\mathbb{R}$, $f$ can be shown to be star-convex under mild conditions of $h$ and $g$.
@@
This structure is common in machine learning optimization tasks (Chen et al., 2021; Jiang et al., 2022), where the function $f$ itself may be nonconvex and nonsmooth. Star-convexity suffices to ensure that Theorem 4.1 remains valid. The only modification is that we need to use the subgradient of $f$ in $\gamma$-PolyakGD (Algorithm 1).
@@
Note that Theorem 4.1 is not applicable to the case $\gamma=2$. To establish a convergence guarantee for this case, we introduce the following inequality. This inequality is an extension of inequality (2.1.10) [^5] This inequality is also known as cocoercivity in the literature (see, e.g., Malitsky and Mishchenko (2024); Latafat et al. (2024)). in Nesterov and others (2018) to Hölder smoothness.
@@
We now consider a more general upper curvature bound than Hölder smoothness, known as the global curvature bound (Nesterov, 2025; Doikov and Nesterov, 2025). This bound is intrinsic as it does not require the objective function $f$ to belong to any specific function class; it is defined by the following quantity:
@@
Given $\hat{\mu}_{f}$, its inverse function $s_{f}$, known as the complexity gauge, corresponds to a stationary criterion for $f$ (see Definition 1 in Nesterov (2025)). Specifically, it satisfies $\hat{\mu}_{f}(s_{f}(t))=t$ for all $t\geq 0$. The following theorem shows that $\gamma$-PolyakGD also adapts to this global curvature bound, matching the result of the universal primal gradient method in Nesterov (2025).
@@
To conclude this section, we extend $\gamma$-PolyakGD to stochastic Polyak stepsize (Loizou et al., 2021). Consider the following stochastic optimization problem
@@ Appendix A Technical Lemmas
According to Theorem 6.3 in Rodomanov and Nesterov (2020), the function $p(x)=\|x\|^{1+\nu}$ is $(\nu,2^{1-\nu}(\nu+1))$-Hölder smooth. Note that $\nabla p_{A}(x)=\nabla(p(Ax))=A\nabla p(Ax)$. Hence, for any $x,y\in\mathbb{R}^{n}$, we have
@@ Appendix B Missing Proofs in Section 3
First, observe that the Huber loss is a quadratic function with condition number $1$ when $\|x\|\leq 1/(2K+1)$. By Theorem 1 in Hazan and Kakade (2019), 1-PolyakGD converges to $x^{\star}$ in one step once the iterate enters this region. Hence, it suffices to consider an initial point $x^{0}$ satisfying $\|x^{0}\|>1/(2K+1)$. We now show that after a single iteration, the iterate $x^{1}$ generated by $1$-PolyakGD satisfies $\|x^{1}\|\leq 1/(2K+1)$. Specifically, the stepsize is given by
@@ Appendix C Missing Proofs in Section 4
Let $\hat{\sigma}_{f}(t)\coloneqq\int_{0}^{t}\frac{\hat{\mu}(\tau)}{\tau}\operatorname{d}\tau$. By Theorem 1 in Nesterov (2025), it holds that
@@
where $(2\hat{\sigma}_{f})^{*}$ is the Fenchel conjugate function of $2\hat{\sigma}_{f}$ (see Definition 3.22 in Drusvyatskiy (2020)). Consequently, by the Fenchel-Young inequality, it follows $\zeta\|\nabla f(x)\|\leq 2\hat{\sigma}_{f}(\zeta)+(2\hat{\sigma}_{f})^{*}(\|\nabla f(x)\|)$. Note that $f(x)-f^{\star}\geq(2\hat{\sigma}_{f})^{*}(\|\nabla f(x)\|)$. We derive
@@
Again, recall that $\zeta$ is arbitrary. This proves Equation (23). Since the function $f$ is convex, by Lemma 4 in Nesterov (2025), it holds that $\hat{\sigma}_{f}(t)\leq\hat{\mu}_{f}(t)$ for any $t\in[0,\infty)$. From Equation (23), we obtain